Using the following returns, calculate the arithmetic average returns, the variances, and the standard deviations for X and Y Returns Year 2 3 4 14% 32 21 24% 45 -22 -25 23 53 Requirement 1 (a) Calculate the arithmetic average return for X. (Click to select) (b) Calculate the arithmetic average return for Y (Click to select) Requirement 2: (a) Calculate the variance for X. (Do not round intermediate calculations.) (Click to select) (b) Calculate the variance for Y. (Do not round intermediate calculations.) (Click to select) Requirement 3 (a) Calculate the standard deviation for X. (Do not round intermediate calculations.) Click to select) (b) Calculate the standard deviation for Y. (Do not round intermediate calculations.) (Click to select)
Answer #1

Assuming the below is population data.

Year

Return X

Return B

(Return X - Average return of X)^2

(Return Y - Average return of Y)^2

1

14.00%

24.00%

0.00160%

0.46240%

2

32.00%

45.00%

3.38560%

7.72840%

3

21.00%

-11.00%

0.54760%

7.95240%

4

-22.00%

-25.00%

12.67360%

17.80840%

5

23.00%

53.00%

0.88360%

12.81640%

Total

68.00%

86.00%

17.49200%

46.76800%

Solution:

As question has not suggested any rounding specification, please use any of the below rounding (six decimal or two decimal):

Six decimal rounding

Two decimal rounding

a

Average Return X = Total X /5 = 68%/5 =

13.600000%

13.60%

b

Average Return Y = Total Y /5 = 86%/5 =

17.200000%

17.20%

a

Variance X = Total of ((Return X - Average return of X)^2)/5 = 17.49200%/5 =

3.498400%

3.4984%

b

Variance Y = Total of ((Return Y - Average return of Y)^2)/5 = 46.76800%/5 =

9.353600%

9.35%

a

Standard deviation X = Variance X^0.5 =

18.704010%

18.70%

b

Standard deviation Y = Variance Y^0.5 =

30.583656%

30.58%